# Personal LLM Agents: Insights and Survey about the Capability, Efficiency and Security

### v1

@@ 5.3 Efficient Memory Manipulation
The Personal LLM Agents need to frequently retrieve internal or external memory to enable more informed decisions. The internal memory is represented as the context tokens and stored as KV cache during the LLM inference. The retrieval of the internal memory is handled implicitly by the self-attention module in Transformer architecture. This leads to LLM conducting more efficient computations over long contexts and trying to minimize the memory footprints while undergoing inference. These issues are similar to the inference efficiency of LLM as discussed in Section 5.1. Therefore, in this subsection, we mainly focus on the efficiency of manipulating external memory, which can be dynamically retrieved and added into the context.
Considering the diverse forms of external memory data, such as user profiles, interaction history, and local raw files (images, videos, etc.), the common practice is to use embedding models [293, 294] to represent memory data with a uniform and high-dimensional vector format. The distance between vectors stands for the semantic similarity between the corresponding data. For each given query, the agent needs to find the most relevant contents in external memory storage. This procedure, together with the maintenance of vectors, can be covered by vector libraries (like Faiss [295, 296, 297] and SCaNN [199]), vector databases [298, 299, 300], or some customized memory structures [301, 302]. Regardless of the functional differences between these systems, their efficiency optimizations are basically targeted at two key aspects, searching and indexing.
5.3.1 Searching Efficiency
@@
The efficiency analysis and optimization can be viewed from different aspects of vector searching. Some aspects are related to search mechanism designs, such as similarity measurement, searching scope, as well as query types, selection, and optimizations. Some aspects, on the other hand, focus on the execution of the search process.
@@
Search Mechanism Design. Multiple similarity criteria can be employed to evaluate the relevance of vector embeddings to a search query, including Hamming Distance, Cosine Distance, and Aggregate Scores [256]. However, the selection of scoring mechanisms lacks stringent principles and often relies on empirical rules [299]. Regarding the types of searches, both approximate and exact $k(\geq 1)$ nearest neighbors [303] search, as well as distance range search, can be utilized to retrieve corresponding vectors. Typically, there exists a trade-off between search accuracy and speed, wherein the use of approximate similarity search or larger k values speeds up the search process but may result in undesirable outcomes. To optimize search latency, rule-based [253, 254] or estimated-cost-based methods [255, 256] are often employed to determine the optimal search plan. These rules and cost models are typically configured offline to avoid unnecessary or time-consuming search actions. To further optimize the search process, hybrid operations that combine vector search with metadata filters are gaining popularity. This involves techniques such as pre-filtering [255, 256, 304], post-filtering, and single-stage filtering [257] to narrow the scope of vector searching. Similarly, the trade-off between search accuracy and speed still exists. For instance, in pre-filtering, metadata analysis is performed before searching, which may lead to overlooking some vectors that do not match the metadata criteria but could be potentially crucial for Personal LLM Agents. Moreover, the utilization of metadata filtering can impede the execution of the search, as it introduces additional computational costs.
@@
5.3.2 Index Optimization
Indexing is also very important to optimize the efficiency of external memory management and retrieval. While comparing the similarity between query vector $q$ and vectors in external memory, a brute-force approach results in a computational complexity of $O(DN)$. However, this approach becomes impractical for scenarios with large vector dimensions ($D$) and dataset sizes ($N$). To address this problem, vector indexing is commonly employed to expedite query searching by reducing the number of required comparisons.
Hardware-aware Index Optimization. Since improving the scalability and efficiency of indexing has become a critical concern, research efforts have focused on hardware-aware approaches to extend external memory capacity while maintaining low latency and high throughput. This is achieved through the utilization of disk-based indexes or the co-design of hardware and algorithms [303]. For example, DiskANN [267] addresses cost-effectiveness by employing a hybrid DRAM-SSD approach. It incorporates Vamana graph indexing on SSDs and employs compressed point representation in DRAM. This configuration enables accurate query responses with less than 10ms latency, even when dealing with a billion-point database. DiskANN++ [268] further improves efficiency by introducing dynamic entry vertex selection and optimizing SSD layout. This enhancement results in a 1.5x to 2.2x increase in Query Per Second (QPS) while maintaining accuracy on real-world datasets. Moreover, CXL-ANNS [269] introduces a collaborative software-hardware approach for scalable approximate nearest neighbor search (ANNS). By utilizing Compute Express Link (CXL), CXL-ANNS disentangles DRAM from the host and consolidates essential datasets into its memory pool. FANNS [270] is a vector search framework on FPGAs, featuring automatic co-design of hardware and algorithms based on user-defined recall requirements and hardware constraints. It supports scale-out with a hardware TCP/IP stack and exhibits notable speedups compared to FPGA and CPU baselines.
@@ 6.1.1 Local Processing
A simple and effective approach to protect user privacy is to perform the computations locally on the users’ personal devices. While LLM service providers are currently working towards improving security and building user trust, it is important to acknowledge that transmitting private data to the cloud inherently introduces additional potential risks. Therefore, processing all data locally is considered a more secure method of interacting with LLMs compared to transmitting data to the cloud. However, deploying LLMs locally poses challenges in efficiently processing user requests due to resource constraints on personal devices. This can lead to slow inference speed or even the inability to perform inference due to the limitations of available memory. Since the data in Personal LLM Agents is mainly processed by the LLM, the key to achieve local computation is to run the LLM on users’ own devices. There are various existing lightweight models [308, 250] and deployment frameworks [309, 274] available for deploying models on edge devices. Furthermore, various model compression techniques [310, 220, 216] are proposed to reduce the model size to further enable the local deployment.
@@ 6.1.3 Data Masking
One direct approach of data masking is to transform the plaintext inputs by hiding or replacing sensitive content such as account numbers, addresses, and personal names. These types of information are commonly referred to as Personally Identifiable Information (PII). However, accurately defining PII can be challenging due to its obscure boundaries and diverse forms, making it difficult to consistently identify and remove it from the original content. The National Institute of Standards and Technology (NIST) has provided a guide [319] that offers recommendations for safeguarding the confidentiality of PII, which could help manage PII more securely.
@@ 6.1.4 Information Flow Control
The aforementioned techniques primarily pertains to the privacy of model input data, while there may also exist the risks of privacy leakage in the model output. This is because the output of the model may not only returns directly to the user but also be sent to other third-party applications, models, users, or intelligent agents. For instance, when an intelligent agent assists a user in making restaurant reservations, it may take the user’s basic profile and schedule information and feed them into the restaurant reservation software. Similarly, when businesses aim to recommend products to users, they may rely on user preference information retrieved from the output of certain personal agents. This method of obtaining privacy information from the output of LLMs is similar to personal data access interfaces in traditional operating systems, where it is crucial to ensure the control and transparency of privacy data access with permission management systems [324]. Transparency necessitates informing users about access information regarding privacy data, including the accessing entity (who), content (what), time (when), intent (why), access method (how), etc.
@@ 6.2.1 Adversarial Attacks
Malicious attacks primarily achieve their objectives through the specialized customization of the model’s inputs or malicious tampering with the model. A significant category of attacks, known as “adversarial attacks”, causes model inference errors by customizing or tampering with the model’s input data, which was initially discovered in image classification models [326]. This type of attacks can induce serious classification errors by adding imperceptible noise to images. Subsequently, researchers have extended this attack method to text data, graph data, and beyond [327]. Such attacks also persist in large langage models [328], which may also accept input of images [329], text [330], and other modalities of data [331] from third parties. For example, when assisting users in automating tasks, attackers may misguide the agent to delete calendar events and leak private conversation data [332], because LLMs often need to input the content of the application’s internal information to generate the next interaction decision. In such cases, if the third-party application feed the LLM with maliciously customized content, it could drive the intelligent agent to engage in unsafe interaction. Traditional defense methods against such attacks in deep learning models usually encompass adversarial defense, abnormal input detection, input preprocessing, output security verification, and more [327]. While these methods theoretically remain applicable to LLM and LLM agents, the large scale of parameters and the characteristics of autoregressive generation may render some computationally expensive methods (such as formalized output security validation and detection of anomalous data based on intermediate layer activations) challenging to implement. Furthermore, some defense methods may require adjustments in the context of LLM. For instance, training the LLM may incur substantial costs, making it impractical to enhance security through adversarial training. Therefore, exploring how to achieve good effects of adversarial defense through parameter-efficient fine-tuning is worth investigating.
@@ 6.2.2 Backdoor Attacks
Another common form of attack is the backdoor attack. Traditional model backdoor attacks are often achieved through data poisoning [333], i.e., inserting maliciously modified samples into the model’s training data, enabling the model to learn deliberate hidden decision logic, such as “when seeing an apple pattern, the model outputs an incorrect classification”. For LLMs, data poisoning may be more challenging due to the huge amount and strict unified management of training data, but another type of backdoor attack methods [334] is still valid, which implants insecure logic into the model by modifying the model input during the test time. Kandpal et al. [335] elicits targeted misclassification when the language models are prompted to perform a particular target task. ProAttack [336] directly utilizes prompts as triggers to inject backdoors into LLMs, which is the first attempt to explore clean-label textual backdoor attacks based on the prompt. PoisonPrompt [337] is a bi-level optimization-based prompt backdoor attack on soft and hard prompt-based LLMs. Since LLMs often use several fixed prompts in certain scenarios, this form of attack, achieved by modifying the prompts, essentially fine-tunes the model’s parameters and thus alters its decision logic. Sun et al. [338] proposed that testing the backward probability of generating sources given targets yields effective defense performance against different types of attacks. Indeed, when attackers mimic normal behavior, this defense method may become ineffective. Therefore, there isn’t a robust solution for backdoor defense in agent systems yet [339]. This highlights the request of developing effective defenses against sophisticated attacks that mimic legitimate behavior.
@@ 6.3 Reliability
In Personal LLM Agents, numerous critical actions are determined by the LLMs, including some sensitive operations such as modifying and deleting user information, procuring services, and sending messages. Therefore, ensuring the reliability of the agent’s decision-making process is crucial. We discuss the reliability of LLMs from three perspectives, including the problems (i.e., where does reliability issues of LLMs manifest from?), improvement (i.e., how can we make the LLMs’ response more reliable?), and inspection (i.e., how can we deal with the LLM’s potentially unreliable output?).
@@ 6.3.1 Problems
Hallucination. LLMs may produce incorrect answers, which can lead to severe consequences. In comparison to LLM-based chatbots that directly interact with users via text, Personal LLM Agents minimize user disruptions by avoiding frequent result verifications, hence amplifying the severity of producing incorrect answers. Researchers have uncovered cases where LLMs generate text that is coherent and fluent but ultimately erroneous. This phenomenon, known as hallucination in natural language processing tasks, poses a challenge to the personal agents as well. Ji et al. [346] delves deeply into the various manifestations of hallucinations in natural language processing tasks. Rawte et al. [347] further discusses the hallucinations in multimodal foundation models, providing valuable references for interested readers.
@@ 6.3.2 Improvement
Alignment. As LLMs grow in size and complexity, concerns have arisen regarding their potential to generate biased, harmful, or inappropriate content. Alignment methods seek to mitigate these risks and ensure that the behavior of LLMs aligns with ethical and societal norms. One common alignment method is the use of pre-training and fine-tuning [352, 353, 354]. LLMs are pre-trained on vast amounts of text data to learn language patterns and representations. During the fine-tuning phase, the models are further trained on more specific and carefully curated datasets, including human-generated examples and demonstrations. This process helps align the models with desired behaviors by incorporating human values and intentions into their training. Another alignment method is reward modeling, which involves defining and optimizing a reward function that reflects desired outcomes or behaviors. By providing explicit rewards or penalties for specific actions, LLMs can be trained to generate outputs that align with those predefined objectives. Reinforcement learning techniques (e.g., RLHF [43], RLAIF [355], C-RLFT [356]) can be employed to optimize the model’s behavior based on these reward signals. oversight and intervention are critical alignment methods. Human reviewers or moderators play a crucial role in reviewing and filtering the outputs of LLMs for potential biases, harmful content, or inappropriate behavior. Their feedback and interventions are used to iteratively improve the model’s performance and align it with desired standards.
@@
Self-Reflection. It has been shown that language models can provide probabilities of providing correct answers [357]. Inspired by the autonomous operation of LLMs, researchers have suggested leveraging the model’s self-reflection to mitigate the problem of incorrect content generation. Huang et al. [211] and Madaan et al. [358] show that LLMs are capable of self-improving with unlabeled data, Shinn et al. [359] propose Reflexion to let LLMs update through its linguistic feedback. Chen et al. [360] propose Self-Debug to iteratively improve the responses on several code generation tasks. SelfCheckGPT [361] allows large models to provide answers to the same input question multiple times and checks the consistency between these responses. If there are contradictions among the answers, there is a higher probability that the model has generated unreliable content. Du et al. [362] attempts to enhance the reliability of model outputs by enabling multiple large model agents to engage in mutual discussion and verification. There are various ways to combine models, similar to the diverse collaboration methods in the human world. However, just as more employees require increased expenses, having more models entails greater computational power requirements. The above works demonstrate a trend in which LLMs are evolving from mere textual generators to intelligent agents, transitioning from primitive comprehension-based reasoning to reflective reasoning with iterative updates.
@@
Retrieval Augmentation. LLMs show strong performance across various tasks, however, the parametric knowledge stored in the models could still be incomplete and difficult to update efficiently. Alternatively, the retrieval augmented methods [199, 200, 363] provide a semi-parametric way to offer complementary non-parametric information, allowing the LLMs to draw on retrieved real-world knowledge when generating content, such as Wikipedia, documents, or knowledge graphs [364]. This approach offers the advantage of not requiring model modification, facilitates real-time information updates, and allows the traceability of generated results to the original data, thereby enhancing the interpretability of the generated information. Retrieval augmentation has been shown effective for traditional pre-trained models such as BERT [365]. However, as for LLMs that already have strong reasoning ability, augmenting the context could also have a negative impact due to irrelevant or noisy information [366]. To tackle these issues, Guo et al. [192] propose a prompt-guided retrieval method for non-knowledge-intensive tasks, enhancing the relevance of retrieved passages for more general queries. Yu et al. [367] propose Chain-of-Note to improve the robustness when dealing with noisy and irrelevant documents. Asai et al. [368] propose Self-RAG to enhance factuality through self-reflection. Wang et al. [369] propose SKR, a self-knowledge-guided retrieval method to balance external knowledge with internal knowledge. Wang et al. [370] propose FLICO to filter the context in advance and improve the fine-grained relevance of retrieved segments. The CRITIC [371] framework utilizes LLMs to verify and iteratively self-correct their output through interaction with external tools, such as a calculator, Python interpreter, and Wikipedia. However, this approach has limited assistance for user requests for which matching content cannot be found in external knowledge bases.

### v2

@@ 5.3 Efficient Memory Manipulation
The Personal LLM Agents need to frequently retrieve external memory to enable more informed decisions, which can depend on the prevailing mechanism called Retrieval-Augmented Generation(RAG). Considering the diverse forms of external memory data, such as user profiles, interaction history, and local raw files (images, videos, etc.), the common practice is to use embedding models [342, 343] to represent memory data with a uniform and high-dimensional vector format. The distance between vectors stands for the semantic similarity between the corresponding data. For each given query, the Personal LLM Agents need to find the most relevant content in external memory storage. The retrieval knowledge then will be injected into Personal LLM Agents through either prompt concatenation or intermediate layer cross-attention [301], with both ways complicating the context of LLM inference. This leads to LLM conducting more efficient computations over long contexts and trying to minimize the memory footprints while undergoing inference, which are similar to improving inference efficiency of LLM as discussed in Section 5.1.
Therefore in this subsection, we mainly focus on the efficient external memory retrieval, which can be considered from two aspects: efficient search and efficient workflow. Efficient search focuses on vector indexing and fast search inside structures like vector libraries (like Faiss [344, 345, 346] and SCaNN [229]), vector databases [347, 348, 349], or some customized memory structures [350, 351] where external memory is stored. While efficient workflow targets to further optimize the end-to-end efficiency of retrieval augmented LLM inference.
5.3.1 Search Efficiency
When comparing the similarity between query vector $q$ and vectors in external memory, a brute-force approach results in a computational complexity of $O(DN)$. However, this approach becomes impractical for scenarios with large vector dimensions ($D$) and dataset sizes ($N$). To alleviate the searching overhead, indexing is commonly employed to expedite query searching by reducing the number of required comparisons.
Hardware-aware Index Optimization. Since improving the scalability and efficiency of indexing has become a critical concern, research efforts have also focused on hardware-aware approaches to extend external memory capacity while maintaining low latency and high throughput. This is achieved through the utilization of disk-based indexes or the co-design of hardware and algorithms [355]. For example, DiskANN [290] addresses cost-effectiveness by employing a hybrid DRAM-SSD approach. It incorporates Vamana graph indexing on SSDs and employs compressed point representation in DRAM. This configuration enables accurate query responses with less than 10ms latency, even when dealing with a billion-point database. DiskANN++ [356] further improves efficiency by introducing dynamic entry vertex selection and optimizing SSD layout. This enhancement results in a 1.5x to 2.2x increase in Query Per Second (QPS) while maintaining accuracy on real-world datasets. Moreover, CXL-ANNS [291] introduces a collaborative software-hardware approach for scalable approximate nearest neighbor search (ANNS). By utilizing Compute Express Link (CXL), CXL-ANNS disentangles DRAM from the host and consolidates essential datasets into its memory pool. FANNS [292] is a vector search framework on FPGAs, featuring automatic co-design of hardware and algorithms based on user-defined recall requirements and hardware constraints. It supports scale-out with a hardware TCP/IP stack and exhibits notable speedups compared to FPGA and CPU baselines.
@@ 5.3.1 Search Efficiency
In terms of the efficiency analysis and optimization of searching itself, some aspects are related to search mechanism design, such as similarity measurement, searching scope, as well as query types, selection, and optimizations. While some aspects, on the other hand, focus on efficient execution of the search process.
@@
Search Mechanism Design. Multiple similarity criteria can be employed to evaluate vector similarity, including Hamming Distance, Cosine Distance, and Aggregate Scores [296]. However, the selection of scoring mechanisms lacks stringent principles and often relies on empirical rules [348]. Regarding the types of searches, both approximate and exact $k(\geq 1)$ nearest neighbors [355] search, as well as distance range search, can be utilized to retrieve corresponding vectors. To optimize search latency, rule-based [293, 294] or estimated-cost-based methods [295, 296] are often employed to determine the optimal search plan. These rules and cost models are typically configured offline to avoid unnecessary or time-consuming search actions. To further optimize the search process, hybrid operations that combine vector search with metadata filters are gaining popularity. This involves techniques such as pre-filtering [295, 296, 354], post-filtering, and single-stage filtering [297] to narrow the scope of vector searching.
@@
5.3.2 Workflow Optimization
No matter for one-shot or iterative RAG system, traditional workflow is sequential, with inference/retrieval stage idle while conducting retrieving/generation. This feature ignores chances of optimization from the potential of execuation parallelism and retrieval locality of requests. Recent studies are working on pipeline and cache techniques to further improve the efficiency of RAG systems.
Pipelining. RaLMSpec [301] is the first work to leverage the advantage of pipeline by enabling a local cache for speculative retrieval. To maintain correctness, a batched verification step is used to guarantee correctness. Besides, cache prefetching, optimal speculation stride scheduler, and asynchronous verification are adopted to further boost the speculation performance. PipeRAG [302] also uses pipeline, and enhance its performance with two different solutions: flexible retrieval intervals and a performance model informed to dynamically adjust the vector search space depending on the latency expectation of the upcoming token in LLM inferences in the pipeline. PipeRAG utilizes an algorithm-system co-design to avoid increasing end-to-end generation latency while optimizing the search quality.
Caching. The reason of selecting cache method arises from the temporal and spatial locality of retrieved documents during different requests, which RaLMSpec [301] has already been utilizing. RAGCache [303] further uses knowledge tree to organize the intermediate states of the retrieved documents both in the GPU and host memory hierarchy. It also presents a prefix-aware Greedy-Dual-Size-Frequency (PGDSF) replacement policy and a cache-aware request scheduling approach to minimize the cache miss rate. Another work, GRITLM [304], trains LM to handle both generative and embedding tasks by distinguishing between them through instructions. Since the common scenario in RAG is an embedding model used for providing relevant context to the generative model to answer user queries, with GRITLM, the embedding and generative model are equivalent, allowing us to conduct Query Caching or Query-Doc Caching and save computation overhead.
@@ 6.1.1 Local Processing
A simple and effective approach to protect user privacy is to perform the computations locally on the users’ personal devices. While LLM service providers are currently working towards improving security and building user trust, it is important to acknowledge that transmitting private data to the cloud inherently introduces additional potential risks. Therefore, processing all data locally is considered a more secure method of interacting with LLMs compared to transmitting data to the cloud. However, deploying LLMs locally poses challenges in efficiently processing user requests due to resource constraints on personal devices. This can lead to slow inference speed or even the inability to perform inference due to the limitations of available memory. Since the data in Personal LLM Agents is mainly processed by the LLM, the key to achieve local computation is to run the LLM on users’ own devices. There are various existing lightweight models [360, 283] and deployment frameworks [361, 308, 362] available for deploying models on edge devices. Furthermore, various model compression techniques [363, 250, 246] are proposed to reduce the model size to further enable the local deployment as discussed in section 5.1.3.
@@ 6.1.2 Secure Remote Processing
Beside HE, Multi-Party Communicatio (MPC) [370] is an important part of traditional applied cryptography, which refer to communication processes involving multiple parties, where several participants need to communicate in a untrusted environment. The challenge of applying MPC in LLM lies in the high computation cost and the significant transition from the mathematical theory of MPC to the actual implementation on LLM. Crypten [371] is a framework that includes common MPC methods, supports standard PyTorch tensor operations, and enables GPU computations.
@@ 6.1.3 Data Masking
One direct approach of data masking is to transform the plaintext inputs by hiding or replacing sensitive content such as account numbers, addresses, and personal names. These types of information are commonly referred to as Personally Identifiable Information (PII). However, accurately defining PII can be challenging due to its obscure boundaries and diverse forms, making it difficult to consistently identify and remove it from the original content. The National Institute of Standards and Technology (NIST) has provided a guide [374] that offers recommendations for safeguarding the confidentiality of PII, which could help manage PII more securely. EmojiCrypt[375] suggested to use emoji to replace user sensitive information, and then use the modified sentences for generation.
@@ 6.1.4 Information Flow Control
The aforementioned techniques primarily pertains to the privacy of model input data, while there may also exist the risks of privacy leakage in the model output. This is because the output of the model may not only returns directly to the user but also be sent to other third-party applications, models, users, or intelligent agents. For instance, when an intelligent agent assists a user in making restaurant reservations, it may take the user’s basic profile and schedule information and feed them into the restaurant reservation software. Similarly, when businesses aim to recommend products to users, they may rely on user preference information retrieved from the output of certain personal agents. This method of obtaining privacy information from the output of LLMs is similar to personal data access interfaces in traditional operating systems, where it is crucial to ensure the control and transparency of privacy data access with permission management systems [380]. Transparency necessitates informing users about access information regarding privacy data, including the accessing entity (who), content (what), time (when), intent (why), access method (how), etc. Evertz et al. [381] proposed a way of evaluating privacy leaks in LLM-integrated systems.
@@ 6.2.1 Adversarial Attacks
Malicious attacks primarily achieve their objectives through the specialized customization of the model’s inputs or malicious tampering with the model. A significant category of attacks, known as “adversarial attacks”, causes model inference errors by customizing or tampering with the model’s input data, which was initially discovered in image classification models [383]. This type of attacks can induce serious classification errors by adding imperceptible noise to images. Subsequently, researchers have extended this attack method to text data, graph data, and beyond [384]. Such attacks also persist in large langage models [385], which may also accept input of images [386], text [387], and other modalities of data [388] from third parties. For example, when assisting users in automating tasks, attackers may misguide the agent to delete calendar events and leak private conversation data [389], because LLMs often need to input the content of the application’s internal information to generate the next interaction decision. In such cases, if the third-party application feed the LLM with maliciously customized content, it could drive the intelligent agent to engage in unsafe interaction. Traditional defense methods against such attacks in deep learning models usually encompass adversarial defense, abnormal input detection, input preprocessing, output security verification, and more [384]. While these methods theoretically remain applicable to LLM and LLM agents, the large scale of parameters and the characteristics of autoregressive generation may render some computationally expensive methods (such as formalized output security validation and detection of anomalous data based on intermediate layer activations) challenging to implement. Furthermore, some defense methods may require adjustments in the context of LLM. For instance, training the LLM may incur substantial costs, making it impractical to enhance security through adversarial training. Therefore, exploring how to achieve good effects of adversarial defense through parameter-efficient fine-tuning is worth investigating. Zhu et al. [390] show that current solutions may be too optimistic: defending against these attacks is possible: adversarial attacks generate unlimited but unreadable gibberish prompts, detectable by perplexity-based filters; manual jailbreak attacks craft readable prompts, but their limited number due to the necessity of human creativity allows for easy blocking. Then they introduce AutoDAN, an interpretable, gradient-based adversarial attack that merges the strengths of both attack types. Guided by the dual goals of jailbreak and readability, AutoDAN optimizes and generates tokens one by one from left to right, resulting in readable prompts that bypass perplexity filters while maintaining high attack success rates, which offers a new way to red-team LLMs and understand jailbreak mechanisms via interpretability.
@@ 6.2.2 Backdoor Attacks
Another common form of attack is the backdoor attack. Traditional model backdoor attacks are often achieved through data poisoning [391], i.e., inserting maliciously modified samples into the model’s training data, enabling the model to learn deliberate hidden decision logic, such as “when seeing an apple pattern, the model outputs an incorrect classification”. For LLMs, data poisoning may be more challenging due to the huge amount and strict unified management of training data, but another type of backdoor attack methods [392] is still valid, which implants insecure logic into the model by modifying the model input during the test time. Kandpal et al. [393] elicits targeted misclassification when the language models are prompted to perform a particular target task. ProAttack [394] directly utilizes prompts as triggers to inject backdoors into LLMs, which is the first attempt to explore clean-label textual backdoor attacks based on the prompt. PoisonPrompt [395] is a bi-level optimization-based prompt backdoor attack on soft and hard prompt-based LLMs. Since LLMs often use several fixed prompts in certain scenarios, this form of attack, achieved by modifying the prompts, essentially fine-tunes the model’s parameters and thus alters its decision logic. Han et al. [396] distill benign knowledge from poisoned pre-trained encoders and transfer it to a new encoder, resulting in a clean pre-trained encoder, which may hurt the LLMs’ performance. Sun et al. [397] proposed that testing the backward probability of generating sources given targets yields effective defense performance against different types of attacks. Indeed, when attackers mimic normal behavior, this defense method may become ineffective. Therefore, there isn’t a robust solution for backdoor defense in agent systems yet [398]. This highlights the request of developing effective defenses against sophisticated attacks that mimic legitimate behavior.
@@ 6.3 Reliability
In Personal LLM Agents, the LLMs determine numerous critical actions, including some sensitive operations such as modifying and deleting user information, purchasing services, and sending messages. Therefore, ensuring the reliability of the agent’s decision-making process is crucial. We discuss the reliability of LLMs from three perspectives, including the problems (i.e., where does reliability issues of LLMs manifest from?), improvement (i.e., how can we make the LLMs’ response more reliable?), and inspection (i.e., how can we deal with the LLM’s potentially unreliable output?).
@@ 6.3.1 Problems
Hallucination. LLMs may produce incorrect answers, which can lead to severe consequences. In comparison to LLM-based chatbots that directly interact with users via text, Personal LLM Agents minimize user disruptions by avoiding frequent result verifications, hence amplifying the severity of producing incorrect answers. Researchers have uncovered cases where LLMs generate text that is coherent and fluent but ultimately erroneous. This phenomenon, known as hallucination in natural language processing tasks, poses a challenge to personal agents as well. Ji et al. [405] delves deeply into the various manifestations of hallucinations in natural language processing tasks. Rawte et al. [406] further discusses the hallucinations in multimodal foundation models, providing valuable references for interested readers.
@@ 6.3.2 Improvement
Alignment. As LLMs grow in size and complexity, concerns have arisen regarding their potential to generate biased, harmful, or inappropriate content. Alignment methods seek to mitigate these risks and ensure that the behavior of LLMs aligns with ethical and societal norms. One common alignment method is the use of pre-training and fine-tuning [411, 412, 413]. LLMs are pre-trained on vast amounts of text data to learn language patterns and representations. During the fine-tuning phase, the models are further trained on more specific and carefully curated datasets, including human-generated examples and demonstrations. This process helps align the models with the desired behaviors by incorporating human values and intentions into their training. Another alignment method is reward modeling, which involves defining and optimizing a reward function that reflects the desired outcomes or behaviors. By providing explicit rewards or penalties for specific actions, LLMs can be trained to generate output that align with those predefined objectives. Reinforcement learning techniques (e.g., RLHF [44], RLAIF [414], C-RLFT [415]) can be employed to optimize the model behavior based on these reward signals. oversight and intervention are critical alignment methods. Human reviewers or moderators play a crucial role in reviewing and filtering the outputs of LLMs for potential biases, harmful content, or inappropriate behavior. Their feedback and interventions are used to iteratively improve the model’s performance and align it with desired standards.
@@
Self-Reflection. It has been shown that language models can provide probabilities of providing correct answers [416]. Inspired by the autonomous operation of LLMs, researchers have suggested leveraging the model’s self-reflection to mitigate the problem of incorrect content generation. Huang et al. [241] and Madaan et al. [417] show that LLMs are capable of self-improving with unlabeled data, Shinn et al. [418] propose Reflexion to let LLMs update through its linguistic feedback. Chen et al. [419] propose Self-Debug to iteratively improve the responses on several code generation tasks. SelfCheckGPT [420] allows large models to provide answers to the same input question multiple times and checks the consistency between these responses. If there are contradictions among the answers, there is a higher probability that the model has generated unreliable content. Du et al. [421] attempts to improve the reliability of model outputs by enabling multiple large model agents to engage in mutual discussion and verification. There are various ways to combine models, similar to the diverse collaboration methods in the human world. However, just as more employees require increased expenses, having more models entails greater computational power requirements. The above works demonstrate a trend in which LLMs are evolving from mere textual generators to intelligent agents, transitioning from primitive comprehension-based reasoning to reflective reasoning with iterative updates.
@@
Retrieval Augmentation. LLMs show strong performance across various tasks, however, the parametric knowledge stored in the models could still be incomplete and difficult to update efficiently. Alternatively, retrieval-augmented methods [229, 230, 422] provide a semi-parametric way to offer complementary nonparametric information, allowing LLMs to draw on retrieved real-world knowledge when generating content, such as Wikipedia, documents, or knowledge graphs [423]. This approach offers the advantage of not requiring model modification, facilitates real-time information updates, and allows the traceability of generated results to the original data, thereby enhancing the interpretability of the generated information. Retrieval augmentation has been shown to be effective for traditional pre-trained models such as BERT [424]. However, for LLMs that already have strong reasoning ability, augmenting the context could also have a negative impact due to irrelevant or noisy information [425]. To tackle these issues, Guo et al. [222] propose a prompt-guided retrieval method for non-knowledge-intensive tasks, enhancing the relevance of retrieved passages for more general queries. Yu et al. [426] propose Chain-of-Note to improve the robustness when dealing with noisy and irrelevant documents. Asai et al. [427] propose Self-RAG to enhance factuality through self-reflection. Wang et al. [428] propose SKR, a self-knowledge-guided retrieval method to balance external knowledge with internal knowledge. Wang et al. [429] propose FLICO to filter the context in advance and improve the fine-grained relevance of retrieved segments. The CRITIC [430] framework utilizes LLMs to verify and iteratively self-correct their output through interaction with external tools, such as a calculator, a Python interpreter, and Wikipedia. Zhang et al. [431] propose RAFT, a retrieval augmented fine-tuning for improving domain specific question answering. However, these approaches still rely on high-performance texts retriever and have limited assistance for user requests for which matching content cannot be easily found in external knowledge bases.